Find four consecutive even integers that sum to 1350.
Answer by Alan3354(69443)

Find four consecutive even integers that sum to 1350.
----------
1350/4 = 337.5, the center
334 + 336 + 338 + 340 = 1348
-----
336 + 338 + 340 + 342 = 1356
No such even integers